# This work is supported by the ML4Seismic Industry Partners at Georgia Tech

### v1

@@ Abstract
Machine learning has taken a critical role in seismic interpretation workflows, especially in fault delineation tasks. However, despite the recent proliferation of pretrained models and synthetic datasets, the field still lacks a systematic understanding of the generalizability limits of these models across seismic data representing a variety of geologic, acquisition and processing settings. Distributional shifts between different data sources, limitations in fine-tuning strategies and labeled data accessibility, and inconsistent evaluation protocols all represent major roadblocks in the deployment of reliable and robust models in real-world exploration settings. In this paper, we present the first large-scale benchmarking study explicitly designed to provide answers and guidelines for domain shift strategies in seismic interpretation. Our benchmark encompasses over $200$ models trained and evaluated on three heterogeneous datasets (synthetic and real data) including FaultSeg3D, CRACKS, and Thebe. We systematically assess pretraining, fine-tuning, and joint training strategies under varying degrees of domain shift. Our analysis highlights the fragility of current fine-tuning practices, the emergence of catastrophic forgetting, and the challenges of interpreting performance in a systematic manner. We establish a robust experimental baseline to provide insights into the tradeoffs inherent to current fault delineation workflows, and shed light on directions for developing more generalizable, interpretable and effective machine learning models for seismic interpretation. The insights and analyses reported provide a set of guidelines on the deployment of fault delineation models within seismic interpretation workflows.
@@ I Introduction
Recent advances in machine learning (ML) have brought about a tectonic shift in the seismic interpretation workflow. ML-assisted approaches are leveraged in different parts of the pipeline, specifically in automated fault detection. Faults are geologically critical features that control fluid flow in Earth’s subsurface, influence reservoir compartmentalization, and pose drilling hazards in hydrocarbon exploration [1, 2]. Beyond the energy sector, faults play a central role in earthquake nucleation and propagation, as well as geohazard and risk assessment in tectonically active regions [3, 4]. Accurate and scalable fault interpretation is therefore a high-impact task across several geoscience domains. Early ML work in this area often repurposed semantic segmentation models from computer vision (such as those developed for natural images) to detect discontinuities in seismic sections. However, these generic frameworks struggled with the unique characteristics of seismic data, particularly the thin, curvilinear geometry of faults and the presence of acquisition artifacts and structural noise. As a result, the field has increasingly shifted toward fault delineation, a task-specific variant of segmentation that emphasizes the extraction of coherent fault structures rather than general pixel-wise classification. This shift has prompted the development of domain-specific feature encoders, adapted loss functions, and structural priors that better capture the morphological signatures of faults [5, 6]. In general, the growing availability of ML methods has accelerated the adoption of data-driven interpretations that have provided new insights into fault structures within large seismic volumes [5, 6].
Figure 1: Typical ML-assisted seismic interpretation workflow
@@
An ML-assisted seismic interpretation pipeline is illustrated in Fig. 1. In a standard workflow, interpreters at a new site begin by leveraging existing labeled datasets or synthetically generated volumes to train machine learning models. These models capture broad geophysical patterns and serve as a foundational prior, which can then be adapted to the new site through fine-tuning on a smaller set of locally interpreted seismic data. After fine-tuning, the models are evaluated within a consistent framework to ensure that their outputs align with geological expectations and interpretation standards. The validated predictions subsequently guide subsurface decision-making, and the new volume may in turn contribute to the pool of training data for future sites. However, while conceptually straightforward, this workflow embeds multiple assumptions and design choices that present significant challenges in practice, which we highlight in light blue in the diagram.
(a) Distributional Shift
(b) Domain-aware Fine-tuning
(c) Standardized Evaluation
Figure 2: Existing challenges in the current interpretation workflow
@@
We further illustrate these challenges in Fig.  2: two central and intertwined challenges in the interpretation pipeline arise from the diversity of data sources (Fig.  2 (a)) and the strategies used to adapt models across them (Fig. 2 (b)). Seismic datasets differ widely in their geological characteristics, acquisition parameters, and resolution, among other factors. Synthetic data, while providing a fully specified ground truth, often fails to capture the complexity and variability of field data. Prior datasets from other regions, though more realistic, may still differ significantly from new target volumes. These domain shifts, whether from synthetic-to-real or across real-world basins, can cause pretrained models to generalize poorly to new sites [7], and eventually provide weak performance for the interpretation task.
@@
Fig. 2(b) showcases a simple example of these aforementioned issues. The figure contains two rows, each displaying three seismic sections. The first row presents the output of three segmentation models commonly used in the seismic community, trained under the same settings with the same training data. We can observe that the three predictions differ significantly from each other. In the second row of Fig. 2(b), three segmentation outputs are shown after applying three different fine-tuning strategies. The first output in the second row is obtained by a model trained from scratch on the target data, while the other two are produced by models pretrained on other (larger) datasets and then finetuned on the target data. We further conduct objective evaluation of these outputs and summarize these metrics in Fig. 3. The y-axis in this figure is the average Hausdorff distance between the output of the model and the ground truth, where smaller values signify better predictions. There are four strategies shown. The first one is where the model is trained on the target volume. The second strategy is to use a model that was trained on a real dataset and fine-tune it on the target dataset. The third strategy is to use a model that was trained in a synthetic dataset and then fine-tuned on the target dataset. These three strategies match the three depicted in Fig. 2(b). A fourth strategy is to use a model that was jointly trained on both synthetic and real data, then fine-tuned on the target data. Although we will discuss these experiments in more detail later in the paper, the figure here shows the large difference in performance across these four strategies. For now, this shows that despite this growing body of work, the field still lacks a systematic understanding of when, how, and why a given strategy outperforms others under domain shift. This paper is the first attempt, to our knowledge, within the community to provide answers and guidelines for domain shift strategies in seismic interpretation.
@@
The third critical complication in the pipeline arises in evaluation, as depicted in Fig. 2(c). Fault labels are derived from expert interpretation and are often subjective [24, 25], particularly in complex or ambiguous structural settings. This subjectivity leads to inconsistencies in ground truth across datasets or among annotators, and complicates objective model evaluation. Furthermore, a given metric may penalize correct but unannotated predictions, while failing to reflect geological plausibility or structural consistency [26, 5]. To address this, our study provides an in-depth analysis of commonly used evaluation metrics (such as Dice coefficient, Chamfer and Hausdorff distances) in the context of fault interpretation. This analysis highlights the need for a more holistic evaluation approach that integrates quantitative measures with geological interpretability.
@@
This paper presents a large-scale benchmarking study regarding the performance variability of over $200$ fault delineation models across training regimes, data characteristics, and architectural choices. These fault delineation models include a combination of eight deep learning architectures, trained and fine-tuned on three distinct datasets, including both synthetic and real volumes. Our experiments cover a broad spectrum of training strategies, including pretraining on a single fault dataset, pre-training on large-scale natural image dataset such as ImageNet [27], using randomly initialized models, and jointly training on multiple fault datasets. We shed light on the impact of these training strategies, with a focus on three critical challenges in the seismic machine learning domain: model generalizability under domain shift, training dynamics, and evaluation robustness. The results of our extensive study allow the community to examine the limitations of the current seismic interpretation pipelines, and underscore the need for a framework that integrates both quantitative metrics and qualitative expert review. We identify failure cases and inconsistencies in fault delineation tasks, and thus highlight the need for interpretability and domain-awareness in model design and evaluation. We provide a succinct highlight of the key observations of our benchmarking experiments in Table II.
@@ II Related Work
The structure of this section mirrors the stages outlined in Fig. 1. Existing literature on machine learning for fault delineation follows a similar modular approach: beginning with data preprocessing and input normalization, proceeding through model training, adaptation strategies and inferential behavior, and culminating in evaluation protocols. In this section, we review prior work organized along these stages, highlighting how each component contributes to generalization across seismic domains.
@@ II-A Data Preparation and Preprocessing
Since raw seismic amplitudes can span orders of magnitude and contain acquisition artifacts, it is a standard practice to normalize each volume either via min–max scaling or z-score transformation. For min–max, each amplitude $v_{ij}$ becomes
$v^{\prime}_{ij}=2\,\frac{v_{ij}-\lambda_{\min}}{\lambda_{\max}-\lambda_{\min}}-1,$ (1)
where $\lambda_{\min}$ and $\lambda_{\max}$ are the global minimum and maximum amplitudes in the volume. Alternatively, z-score normalization
$v^{\prime\prime}_{ij}=\frac{v_{ij}-\mu}{\sigma}$ (2)
centers the data to zero mean and unit variance, with $\mu$ and $\sigma$ computed over all voxels. Fault masks (either interpreter‐drawn or synthetically generated) are saved as binary images (.png or .npy) and cropped identically to their corresponding seismic windows. This systematic preprocessing pipeline (from raw SEG-Y/DAT ingestion, through standardized normalization and windowed slicing, to precisely aligned input–mask pairs) provides a consistent basis for benchmarking and comparing fault delineation models across diverse seismic domains.
@@ II-B Training Dynamics and Setups
Popular network choices designed for natural image segmentation are sensitive to training and design choices like model architecture, input window size, and loss functions. Networks such as U-Net [31], DeepLab [29], and SegFormer [39] typically accept fixed‐size patches (e.g., $256\times 256$ or $512\times 512$ pixels), balancing the need for sufficient context against GPU memory limits. During each training iteration, these patches are sampled, sometimes at multiple scales to expose the network to varied object sizes and to regularize against overfitting. Segmentation models optimize losses designed to reconcile pixel‐wise accuracy with region‐level coherence. For binary masks, a commonly used loss is the pixel‐wise binary cross‐entropy (BCE) [40], given as follows:
$\mathcal{L}_{\mathrm{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}\bigl{[}y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\bigr{]},$ (3)
where $p_{i}$ is the predicted probability at pixel $i$, $y_{i}\in\{0,1\}$ is the ground‐truth label, and $N$ is the number of pixels in the patch. An alternative that tends to better adapt to settings with class imbalance and emphasize overlap is the Dice loss [41] given as follows:
$\mathcal{L}_{\mathrm{Dice}}=1-\frac{2\sum_{i=1}^{N}p_{i}\,y_{i}+\epsilon}{\sum_{i=1}^{N}p_{i}+\sum_{i=1}^{N}y_{i}+\epsilon},$ (4)
where $\epsilon$ is a small constant to provide numerical stability. The BCE loss emphasizes the pixel fidelity while the Dice loss takes the overlap into consideration. Therefore, hybrid objectives combining BCE and Dice, e.g. $\mathcal{L}=\alpha\mathcal{L}_{\mathrm{BCE}}+(1-\alpha)\mathcal{L}_{\mathrm{Dice}}$, are commonplace to leverage both pixel fidelity and region overlap [42, 43].
@@ II-C Generalization and Transferability
Generalization refers to a model’s ability to maintain performance when presented with new, unseen data that (ideally) follows the same distribution as the training set [51, 52, 53]. In segmentation, this means accurately delineating objects or regions under variations in lighting, scale, or background texture. However, a common obstacle to generalization is domain shift, which occurs when the statistical properties of training and test data differ (such as changes in camera sensors or scene composition) often leading to degraded performance [54, 55, 56, 57, 58]. Domain adaptation encompasses strategies to reduce this gap, for instance by aligning feature distributions between source and target domains [59, 60]. Another common technique to adapt to target domains is fine-tuning: a network pretrained on a large, generic dataset (the source) is adapted to a more specialized task (the target) by retraining some or all layers, thereby leveraging learned representations while adjusting to new data characteristics [61, 62].
In seismic segmentation, generalization to new domains is particularly difficult due to variability in acquisition parameters, stratigraphy, frequency content, and noise levels across surveys. This is true for fault delineation, where subtle and discontinuous features are easily obscured by processing artifacts or geologic heterogeneity [63, 47]. Models trained on one domain (such as synthetic datasets like FaultSeg3D [1]) often fail to transfer effectively to real data from different basins. To mitigate such domain shift, researchers have explored domain adaptation techniques including feature-space alignment [64], adversarial learning [65], and seismic-style transfer [66]. Fault-specific approaches also include synthetic fault injection [67] and contrastive learning tailored to structural continuity [20]. Transfer learning pipelines typically rely on pretrained encoders (either from natural images [68, 69] or large-scale seismic simulations with domain-specific decoders [47, 1]) followed by fine-tuning on limited labeled sections. These strategies must navigate a tradeoff between plasticity and stability: aggressive weight updates are prone to inducing catastrophic forgetting [23], while conservative tuning may fail to capture domain-specific features. As a result, generalization in seismic fault delineation requires careful calibration of both model architecture and adaptation strategy to handle the subtle, spatially sparse structures across diverse geological, acquisition, and imaging settings.

### v2

@@ Abstract
Machine learning has taken a critical role in seismic interpretation workflows, especially in fault delineation tasks. However, despite the recent proliferation of pretrained models and synthetic datasets, the field still lacks a systematic understanding of the generalizability limits of these models across seismic data representing a variety of geologic, acquisition and processing settings. Distributional shifts between different data sources, limitations in fine-tuning strategies and labeled data accessibility, and inconsistent evaluation protocols all represent major roadblocks in the deployment of reliable and robust models in real-world exploration settings. In this paper, we present the first large-scale benchmarking study explicitly designed to provide answers and guidelines for domain shift strategies in seismic interpretation. Our benchmark encompasses over $200$ combinations of model architectures, datasets and training strategies, across three diverse datasets (synthetic and real data) including FaultSeg3D, CRACKS, and Thebe. We systematically assess pretraining, fine-tuning, and joint training strategies under varying degrees of domain shift. Our analysis highlights that commonly used fine-tuning practices can lead to catastrophic forgetting, especially when source and target datasets are distributionally disjoint, and that larger models such as Segformer tend to be more robust to adaptation than smaller architectures. Interestingly, we also find that common domain adaptation methods outperform fine-tuning when the distributional shift is large, yet underperform when source and target domains are similar. Finally, we complement conventional segmentation metrics with a novel analysis based on fault characteristic descriptors, revealing how models absorb structural biases from different training datasets. Overall, we establish a robust experimental baseline to provide insights into the tradeoffs inherent to current fault delineation workflows, and shed light on directions for developing more generalizable, interpretable and effective machine learning models for seismic interpretation. The insights and analyses reported provide a set of guidelines on the deployment of fault delineation models within seismic interpretation workflows.
@@ I Introduction
Recent advances in deep learning (DL) have brought about a tectonic shift in the seismic interpretation workflow [1, 2], mirroring broader trends across other domains like geoscience [3, 4], sustainable energy systems [5, 6, 7] and biomedical applications [8, 9, 10, 11, 12, 13, 14]. DL-assisted approaches are leveraged in different parts of the pipeline, specifically in automated fault detection. Faults are geologically critical features that control fluid flow in Earth’s subsurface, influence reservoir compartmentalization, and pose drilling hazards in hydrocarbon exploration [15, 16]. Beyond the energy sector, faults play a central role in earthquake nucleation and propagation, as well as geohazard and risk assessment in tectonically active regions [17, 18]. Accurate and scalable fault interpretation is therefore a high-impact task across several geoscience domains. Early work in this area often repurposed semantic segmentation models from computer vision (such as those developed for natural images) to detect discontinuities in seismic sections. However, these generic frameworks struggled with the unique characteristics of seismic data, particularly the thin, curvilinear geometry of faults and the presence of acquisition artifacts and structural noise. As a result, the field has increasingly shifted toward fault delineation, a task-specific variant of segmentation that emphasizes the extraction of coherent fault structures rather than general pixel-wise classification. This shift has prompted the development of domain-specific neural networks, feature encoders, adapted loss functions, and structural priors that better capture the morphological signatures of faults [19, 20]. In general, the growing availability of DL methods has accelerated the adoption of data-driven interpretations that have provided new insights into fault structures within large seismic volumes [19, 20].
Figure 1: Typical DL-assisted seismic interpretation workflow
@@
A DL-assisted seismic interpretation pipeline is illustrated in Fig. 1 (top diagram). In a standard workflow, interpreters at a new site begin by leveraging existing labeled datasets or synthetically generated volumes to train machine learning models. These models capture broad geophysical patterns and serve as a foundational prior, which can then be adapted to the new site through fine-tuning on a smaller set of locally interpreted seismic data. After fine-tuning, the models are evaluated within a consistent framework to ensure that their outputs align with geological expectations and interpretation standards. The validated predictions subsequently guide subsurface decision-making, and the new volume may in turn contribute to the pool of training data for future sites. However, while conceptually straightforward, this workflow embeds multiple assumptions and design choices that present significant challenges in practice, which we highlight at the bottom of Figure 1.
@@
Two central and intertwined challenges in the interpretation pipeline arise from the diversity of data sources (Fig.  1, orange box) and the strategies used to adapt models across them (Fig. 1, green box). Seismic datasets differ widely in their geological characteristics, acquisition parameters, and resolution, among other factors. Synthetic data, while providing a fully specified ground truth, often fails to capture the complexity and variability of field data. Prior datasets from other regions, though more realistic, may still differ significantly from new target volumes. These domain shifts, whether from synthetic-to-real or across real-world basins, can cause pretrained models to generalize poorly to new sites [21], and eventually provide weak performance for the interpretation task.
@@
Fig. 1 (green box) showcases a simple example of the aforementioned issues. The figure contains two rows, each displaying three seismic sections. The first row presents the output of three segmentation models commonly used in the seismic community, trained under the same settings with the same training data. We can observe that the three predictions differ significantly from each other. In the second row of the box, three segmentation outputs are shown after applying three different fine-tuning strategies. The first output in the second row is obtained by a model trained from scratch on the target data, while the other two are produced by models pretrained on other (larger) datasets and then finetuned on the target data. We further conduct objective evaluation of these outputs and summarize these metrics in Fig. 2. The y-axis in this figure is the average Hausdorff distance between the output of the model and the ground truth, where smaller values signify better predictions. There are four strategies shown. The first one is where the model is trained on the target volume. The second strategy is to use a model that was trained on a real dataset and fine-tune it on the target dataset. The third strategy is to use a model that was trained in a synthetic dataset and then fine-tuned on the target dataset. These three strategies match the three depicted in Fig. 1(green box). A fourth strategy is to use a model that was jointly trained on both synthetic and real data, then fine-tuned on the target data. Although we will discuss these experiments in more detail later in the paper, the figure here shows the large difference in performance across these four strategies. For now, this shows that despite this growing body of work, the field still lacks a systematic understanding of when, how, and why a given strategy outperforms others under domain shift. This paper is the first attempt, to our knowledge, within the community to provide answers and guidelines for domain shift strategies in seismic interpretation.
@@
The third critical complication in the pipeline arises in evaluation, as depicted in Fig. 1 (purple box). Fault labels are derived from expert interpretation and are often subjective [38, 39], particularly in complex or ambiguous structural settings. This subjectivity leads to inconsistencies in ground truth across datasets or among annotators, and complicates objective model evaluation. Furthermore, a given metric may penalize correct but unannotated predictions, while failing to reflect geological plausibility or structural consistency [40, 19]. To address this, our study provides an in-depth analysis of commonly used evaluation metrics (such as Dice coefficient, Chamfer and Hausdorff distances) in the context of fault interpretation. This analysis highlights the need for a more holistic evaluation approach that integrates quantitative measures with geological interpretability.
@@
This paper presents a large-scale benchmarking study regarding the performance variability of finetuning strategies for fault delineation models across training regimes, data characteristics, and architectural choices. These fault delineation models include a combination of eight deep learning architectures, trained and fine-tuned on three distinct datasets, including both synthetic and real volumes. Our experiments cover a broad spectrum of training strategies, including pretraining on a single fault dataset, pre-training on large-scale natural image dataset such as ImageNet [41], using randomly initialized models, jointly training on multiple fault datasets, and using standard domain adaptation (DA) strategies. Our contributions are as follows:
  • We present the first large-scale benchmarking study focused on fault delineation under domain shift, spanning over 200 experimental setups across 8 DL architectures, 5 pretraining-finetuning datasets and 3 evaluation ones, as well as several training configurations.
  • We systematically evaluate pretraining, fine-tuning, and joint training strategies, exposing failure cases such as catastrophic forgetting and dataset-dependent brittleness.
  • We analyze the effect of architectural scale and design on transferability, highlighting that larger models often adapt more effectively to finetuning, and smaller models benefit from domain adaptation strategies under large domain shift.
  • We compare evaluation metrics and highlight their limitations, advocating for evaluation practices that incorporate structural and geological plausibility. We also introduce a novel analysis based on fault characteristic metrics, which sheds new light on the way data properties and model behavior are entangled, and opens up new avenues from which to understand finetuning and domain transfer strategies.
  • We open-source our code and models to provide a foundational reference to scientists and practitioners.
The remainder of this paper is organized as follows. Section II reviews prior work on machine learning for fault delineation in the context of seismic interpretation workflows. Section III describes the three datasets used in our benchmarking study, highlighting their geological and statistical characteristics. Section IV details our experimental setup, including data preparation, model architectures, and evaluation protocols. Section V presents the results of over 200 training configurations, analyzed across generalization and transferability, training dynamics, and metric evaluation. Section VI introduces an additional analysis on geometric and topological metrics to characterize dataset–model interactions. Finally, Section VII concludes the paper by summarizing key findings, outlining practical guidelines, and identifying open challenges for future research.
@@ II Related Work
The structure of this section mirrors the stages outlined in Fig. 1. Existing literature on machine learning for fault delineation follows a similar modular approach: beginning with data preprocessing and input normalization, proceeding through model training, adaptation strategies and inferential behavior, and culminating in evaluation protocols. To provide both breadth and depth, each subsection first summarizes the general deep learning principles that motivate a given stage, and then narrows the discussion to prior work in seismic interpretation and fault segmentation specifically. In this way, we highlight how established DL techniques intersect with geophysical challenges, and how each component contributes to generalization across seismic domains.
@@ II-A Data Preparation and Preprocessing
Since raw seismic amplitudes can span orders of magnitude and contain acquisition artifacts, it is a standard practice to normalize each volume either via min–max scaling or z-score transformation. Fault masks (either interpreter‐drawn or synthetically generated) are saved as binary images (.png or .npy) and cropped identically to their corresponding seismic windows. This systematic preprocessing pipeline (from raw SEG-Y/DAT ingestion, through standardized normalization and windowed slicing, to precisely aligned input–mask pairs) provides a consistent basis for benchmarking and comparing fault delineation models across diverse seismic domains.
@@ II-B Training Dynamics and Setups
Popular network choices designed for natural image segmentation are sensitive to training and design choices like model architecture, input window size, and loss functions. Networks such as U-Net [45], DeepLab [43], and SegFormer [53] typically accept fixed‐size $n\times n$ patches (usually $n\in\{128,256,512\}$), balancing the need for sufficient context against GPU memory limits. During each training iteration, these patches are sampled, sometimes at multiple scales to expose the network to varied object sizes and to regularize against overfitting. Segmentation models optimize losses designed to reconcile pixel‐wise accuracy with region‐level coherence. The binary cross-entropy (BCE) loss [54] focuses on classifying each pixel correctly, while the Dice loss [55] measures the overlap between predicted and ground-truth fault regions, making it well-suited for imbalanced datasets. In practice, many works combine BCE and Dice to take advantage of both pixel fidelity and shape alignment [56, 57].
@@ II-C Generalization and Transferability
Generalization refers to a model’s ability to maintain performance when presented with new, unseen data that (ideally) follows the same distribution as the training set [65, 66, 67]. In segmentation, this means accurately delineating objects or regions under variations in lighting, scale, or background texture. However, a common obstacle to generalization is domain shift, which occurs when the statistical properties of training and test data differ (such as changes in camera sensors or scene composition) often leading to degraded performance [68, 69, 70, 71, 72]. Domain adaptation encompasses strategies to reduce this gap when the target data is limited, for instance by aligning feature distributions between source and target domains [73, 74]. Another common technique to adapt to target domains is fine-tuning: a network pretrained on a large, generic dataset (the source) is adapted to a more specialized task (the target) by retraining some or all layers, thereby leveraging learned representations while adjusting to new data characteristics [75, 76].
In seismic segmentation, generalization to new domains is particularly difficult due to variability in acquisition parameters, stratigraphy, frequency content, and noise levels across surveys. This is especially true for fault delineation, where subtle and discontinuous features are easily obscured by processing artifacts or geologic heterogeneity [77, 61]. Models trained on one domain (such as synthetic datasets like FaultSeg3D [15]) often fail to transfer effectively to real data from different basins. Recent surveys [78, 79] provide comprehensive overviews over the way deep learning approaches in exploration seismology struggle with such domain shifts and emphasize fault segmentation as a particularly brittle task.
To mitigate these challenges, researchers have explored a variety of domain adaptation strategies, including feature-space [80] and frequency-space [81] alignment, adversarial learning [82], and seismic-style transfer [83]. Fault-specific adaptations include synthetic fault injection to enrich training distributions [84], contrastive learning schemes designed for curvilinear structures [34], and self-supervised pretraining frameworks that improve transferability across surveys with limited labels [19, 20]. Transfer learning pipelines typically rely on pretrained encoders (either from natural images [85, 86] or large-scale seismic simulations with domain-specific decoders [61, 15]) followed by fine-tuning on limited labeled sections. These strategies must navigate a tradeoff between plasticity and stability: aggressive weight updates are prone to inducing catastrophic forgetting [37], while conservative tuning may fail to capture domain-specific features. As a result, generalization in seismic fault delineation requires careful calibration of both model architecture and adaptation strategy to handle the subtle, spatially sparse structures across diverse geological, acquisition, and imaging settings.